# Perturbations of relativistic dissipative stars

### v1

@@ 5.2 Viscous Regime
$=$ $16\pi\eta e^{\Phi/2}\frac{\partial\psi}{\partial t}-\frac{8\pi e^{\Phi}}{r}\Bigl{(}2\frac{d\eta}{dr_{\star}}+\eta\frac{d\Phi}{dr_{\star}}\Bigr{)}\beta\,.$
@@
where the coefficients $\mathsf{b}_{1,2,3}$ and $\mathsf{c}_{1,2,3,4}$ are all proportional to viscous transport coefficients, and are written in the SM. Therefore, axial perturbations in the presence of viscosity are described by two coupled wave equations, where both of them include explicitly dissipative terms. This shows that, in the presence of viscosity, there is a new family of modes. These modes do not have a counterpart in the perfect fluid case, since the restoring force is provided solely by shear viscosity. This is one of our main results, which highlights the need of a self-consistent treatment of dissipative effects to study relativistic oscillations of fluids. In the perfect fluid case, the axial sector describes only a propagating GW. Estimates based on energy balance Lindblom and Detweiler (1983) can perhaps guess the dissipation of GWs through absorption by the fluid, but they fail to predict the existence of a fluid mode. Here we have shown that these fluid modes exist, and propagate with a speed characterized by the ratio between the shear viscosity $\eta$ and the heat dissipation timescale $\tau_{\mathcal{Q}}$.
@@ 6.1 Inviscid Regime
One can notice immediately that the equation ~$\tilde{\mathcal{E}}_{S}$ implies that $h_{g}=0$, both in the inviscid and in the viscous case. Therefore we are left with only $6$ variables. Although there are approaches that reduce the problem solely to equations for the metric perturbations Chandrasekhar and Ferrari (1991); Gundlach and Martin-Garcia (2000), here we write down a system of three coupled wave equations for the two metric perturbations $\{h_{p},h_{S}\}$ and the fluid perturbation $\omega$ Allen et al. (1998). Additionally there is a constraint which can be enforced in replacement of the corresponding wave equation, resulting in a more accurate time domain evolution Nagar et al. (2004).
@@
Now, using these, we can use the combination $\mathcal{E}_{g}-\mathcal{E}_{p}$ to obtain the constraint equation
@@
$-\frac{32\pi\rho r^{2}e^{\Lambda}}{c_{s}^{2}}H+e^{\Lambda}\Bigl{(}\lambda^{2}+2-16\pi r^{2}\mathrm{e}\Bigr{)}h_{p}\Biggr{]}\,.$
@@
$+e^{\Phi}\Biggl{[}16\pi\rho\Bigl{(}1-\frac{1}{c_{s}^{2}}\Bigr{)}H+\Biggl{(}\frac{4e^{-\Lambda}}{r^{2}}-\frac{\lambda^{2}+4}{r^{2}}+8\pi(3\mathrm{e}+\mathrm{p})+e^{-\Phi}\Bigl{(}\frac{d\Phi}{dr_{\star}}\Bigr{)}^{2}\Biggr{)}h_{p}\Biggr{]}\,,$
@@
These equations, together with the constraint (42), are equivalent to Equations (9)–(11) in Nagar et al. (2004), where the correspondence between the variables is written in Table 2.
@@
From these equations we can clearly see that $\{h_{S},h_{p}\}$ encode gravitational perturbations, propagating at the local speed of light, whereas $H$ is the fluid perturbation, which propagates at the local sound speed. The equations are coupled in a non–trivial way, but only first order derivatives appear as source terms.
@@ 6.2 Inverse Cowling Approximation
In order to build the physical intuition regarding the impact of viscosity for even parity perturbations, we begin by studying the problem in the ICA Andersson et al. (1996). This approximation is not precisely gauge invariant, and its physical information is limited Wu and Leung (2007). However, it will be useful to unveil some of the strucutre that viscosity adds to the equations. We will freeze the fluid perturbations, and study only the evolution of the gravitational modes. This was used, e.g., to confirm that the $w$–modes found in the polar sector could be thought of precisely as gravitational modes Kokkotas and Schmidt (1999). In this case, freezing the fluid degrees of freedom in the presence of viscosity simplifies remarkably the equations, and unveils partially the role of viscosity in the propagation of gravitational degrees of freedom. After some lengthy algebra one can show that the even parity equations reduce to the following two coupled wave equations
$-\frac{\partial^{2}h_{S}}{\partial t^{2}}+\frac{\partial^{2}h_{S}}{\partial r_{\star}^{2}}=$ $\Bigl{(}\dots\Bigr{)}_{\mathrm{P}.F.}-2\pi e^{\Phi/2}V_{+}\Bigl{(}\frac{\partial h_{p}}{\partial t}+4\frac{\partial h_{S}}{\partial t}-\frac{\partial h_{q}}{\partial r_{\star}}\Bigr{)}+16\pi\eta e^{\Phi/2}\frac{\partial h_{S}}{\partial t}$ (44)
@@
where the $(\dots)_{\mathrm{P}.F.}$ denotes the value of the right hand side of the equivalent equation in the perfect fluid case (i.e., equations (43), setting all the fluid variables to zero). The Hamiltonian constraint is also modified, obtaining
@@
Although not shown here, the equation for the first order time derivative of $h_{q}$ is also modified. In the modified wave equations we can see terms that are directly proportional to first order time derivatives. These are typically associated to dissipative processes. In this case, the dissipative constant is controlled both by $V_{+}=\zeta+4\eta/3+\rho(\tau_{\mathrm{e}}-\tau_{\mathrm{p}})$, as well as the shear viscosity directly, $\eta$. The transport coefficient $\tau_{\mathcal{Q}}$ does not appear altogether. The bulk viscosity only appears through $V_{+}$, and the remaining transport coefficients only appear through $\tau_{-}$. This leads to the expectation that, although the phase space of transport coefficients is large, in practice, it will be dominated by a reduced number of combinations of parameters.
@@ 6.3 Viscous Regime
The remaining equations, on the other hand, are very lengthy and unilluminating. We provide a ready to use Wolfram language package, as well as scripts containing the equations themselves, to make the equations available Redondo-Yuste (2024). We provide these equations in a way that is ready for a numerical implementation, either in the time or in the frequency domain. The only things left to be specified are the dependence of the transport coefficients on the fundamental properties of the star, e.g., $\eta=\eta(\mathrm{e})$, and the equation of state, $\mathrm{p}=\mathrm{p}(\mathrm{e})$.
@@ 7 Conclusions
We have employed the formalism of Gerlach and Sengupta Gerlach and Sengupta (1979a, b, 1980), as applied by Gundlach and Martín-García Martin-Garcia and Gundlach (1999); Gundlach and Martin-Garcia (2000); Martin-Garcia and Gundlach (2001) to reduce the perturbed Einstein equations to a set of scalar equations, in terms of gauge invariant variables. We have focused our attention on radiative multipoles, i.e., multipoles with $\ell\geq 2$, leaving the study of the lower multipoles (including radial perturbations) to future work.
@@
In the polar sector, the situation is more complicated. We have first checked that we recover the perfect fluid results in the appropriate limit. Then, we show how these are modified in the ICA, i.e., freezing the fluid degrees of freedom. In this case, there are two wave equations for gravitational wave modes, as well as one constraint. Remarkably, the equations are modified through the inclusion of dissipative terms, i.e., terms proportional to first order time derivatives, proportionally only to two particular combinations of the transport coefficients. This shows that although the dimensionality of the parameter space of transport coefficient is large, most of the dynamics is perhaps captured by some simple combinations of transport coefficients.
Finally, we provide a Wolfram language package Redondo-Yuste (2024) and scripts with the perturbative equations including all dissipative effects, also in the polar sector. We studied their causal structure. We show the existence of two propagating degrees of freedom associated to GWs (one of them is killed by a constraint), as well as three fluid modes. One of these fluid modes has the same propagation speed as the novel fluid mode in the axial sector, which we dub a shear mode. This mode does not have a perfect fluid counterpart. We argue that the last pair of modes is composed by the usual fluid mode in the even parity sector (associated to density fluctuations), whose propagation speed is modified in the presence of viscosity, and another mode which is purely viscous. Remarkably, we are able to recover the causality constraints of the BDNK theory from the analysis of the principal part of these equations, therefore checking the validity of our results.
We did not study the effect of non radiating multipoles ($\ell=0,1$), as well as the junction conditions that match these perturbations to the perturbations in the exterior of the star (described by the Regge–Wheeler and Zerilli equations, respectively). Moreover, we simplified the problem by assuming that the thermal conductivity was negligible. These are all directions in which we plan to extend our work in the future. More interestingly, the equations provided here allow us to improve our understanding of the interaction between GWs and matter. In a companion work Boyanov et al. (2024) we have shown how this occurs for axial perturbations. Studying scattering properties, as well as the spectrum of compact stars under even parity perturbations, using the equations provided in this work, is a natural continuation of this.
@@ Acknowledgements
I am indebted to the developers and maintainers of xAct, which made possible most of this work Martín-García (2008).
@@
I acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation.
@@ Appendix A Gauge Invariance of the perturbations
where we remind the reader that $D_{a}$ is the covariant derivative on the $2$–sphere, and $\mathcal{Y}^{lm}$ are the spherical harmonics, defined as the eigenfunctions of the equation
@@
An arbitrary vector field can also be decomposed as
@@ Appendix C Coefficients of the odd parity equations
$\mathsf{b}_{1}=$ $\frac{4\pi re^{(\Phi+\Lambda)/2}}{\mathcal{K}}\Biggl{(}2\frac{\partial\log\eta}{\partial r_{\star}}+\frac{\partial\Phi}{\partial r_{\star}}\Biggr{)}\,,\hskip 25.0pt\mathsf{b}_{2}=-e^{\Phi/2}\Bigl{(}\frac{1}{\eta}+16\pi\tau\Bigr{)}\,,$
@@
$\hskip 85.35826pt+re^{(\Phi+\Lambda)/2}\Bigl{(}\frac{\partial\Lambda}{\partial r_{\star}}+3\frac{\partial\Phi}{\partial r_{\star}}\Bigr{)}\Biggr{]}\,,$
@@
$\mathsf{c}_{2}=$ $e^{-\Lambda/2}\Biggl{[}\tau+4\pi r\Biggl{(}re^{\Lambda}\frac{\partial\Phi}{\partial r_{\star}}-6e^{(\Phi+\Lambda)/2}\Biggr{)}\frac{\eta}{\mathcal{K}}-8\pi r^{2}e^{\Lambda}\frac{\eta}{\mathcal{K}}\frac{\partial\log\eta}{\partial r_{\star}}\Biggr{]}\,,$
@@
$\mathcal{K}=\frac{\partial\Lambda}{\partial r_{\star}}+\frac{\partial\Phi}{\partial r_{\star}}=8\pi re^{(\Phi+\Lambda)/2}(\mathrm{e}+\mathrm{p})\,.$ (C.2)
@@
$\frac{d\beta}{dt}=-e^{-\Phi/2}\Biggl{(}r\frac{d\psi}{dr_{\star}}+e^{(\Phi-\Lambda)/2}\psi\Biggr{)}\,,$ (C.3)

### v2

@@ 5.2 Viscous Regime
[…]
@@
where the coefficients $\mathsf{b}_{1,2,3}$ and $\mathsf{c}_{1,2,3,4}$ are all proportional to viscous transport coefficients, and are written in the SM. Therefore, axial perturbations in the presence of viscosity are described by two coupled wave equations, where both of them include explicitly dissipative terms. An analysis of the characteristics of the problem will rapidly show two (pairs of) modes propagating with characteristic velocities $1$ and $\sqrt{\eta\tau_{\mathcal{Q}}/(\mathrm{e}+\mathrm{p})}$, respectively. This shows that, in the presence of viscosity, there is a new family of modes. These modes do not have a counterpart in the perfect fluid case, since the restoring force is provided solely by shear viscosity. This is one of our main results, which highlights the need of a self-consistent treatment of dissipative effects to study relativistic oscillations of fluids. In the perfect fluid case, the axial sector describes only a propagating GW. Estimates based on energy balance Lindblom and Detweiler (1983) can perhaps guess the dissipation of GWs through absorption by the fluid, but they fail to predict the existence of a fluid mode. Here we have shown that these fluid modes exist, and propagate with a speed characterized by the ratio between the shear viscosity $\eta$ and the dissipation timescale $\tau_{\mathcal{Q}}$.
@@ 6.1 Inviscid Regime
One can notice immediately that the equation $\tilde{\mathcal{E}}_{S}$ implies that $h_{g}=0$, both in the inviscid and in the viscous case. Therefore we are left with only $6$ variables. Although there are approaches that reduce the problem solely to equations for the metric perturbations Chandrasekhar and Ferrari (1991); Gundlach and Martin-Garcia (2000), here we write down a system of three coupled wave equations for the two metric perturbations $\{h_{p},h_{S}\}$ and the fluid perturbation $\omega$ Allen et al. (1998). The wave character of the equations is discussed in more detail in 6.3. Additionally there is a constraint which can be enforced in replacement of the corresponding wave equation, resulting in a more accurate time domain evolution Nagar et al. (2004).
@@
Now, using these, we can use the combination $\mathcal{E}_{g}-\mathcal{E}_{p}$ to obtain the equation
@@
[…]
@@
[…]
@@
These equations, together with Eq. (42), are equivalent to Equations (9)–(11) in Nagar et al. (2004), where the correspondence between the variables is written in Table 2.
@@
From these equations we can clearly see that $\{h_{S},h_{p}\}$ encode gravitational perturbations, propagating at the local speed of light, whereas $H$ is the fluid perturbation, which propagates at the local sound speed. We could use Eq. (42) instead of the first equation in Eqs. (43) to evolve $h_{S}$. In that case, there would only remain one propagating degree of freedom for the gravitational sector, as expected Nagar et al. (2004). The equations are coupled in a non–trivial way, but only first order derivatives appear as source terms.
@@ 6.2 Inverse Cowling Approximation
In order to build the physical intuition regarding the impact of viscosity for even parity perturbations, we begin by studying the problem in the ICA Andersson et al. (1996). This approximation is not precisely gauge invariant, and its physical information is limited Wu and Leung (2007). However, it will be useful to unveil some of the strucutre that viscosity adds to the equations. We will freeze the fluid perturbations, and study only the evolution of the gravitational modes. This was used, e.g., to confirm that the ww–modes found in the polar sector could be thought of precisely as gravitational modes Kokkotas and Schmidt (1999). In this case, freezing the fluid degrees of freedom in the presence of viscosity simplifies remarkably the equations, and unveils partially the role of viscosity in the propagation of gravitational degrees of freedom. After some lengthy algebra one can show that the even parity equations reduce to the following two coupled equations
[…]
@@
where the $(\dots)_{\mathrm{P}.F.}$ denotes the value of the right hand side of the equivalent equation in the perfect fluid case (i.e., equations (43), setting all the fluid variables to zero). Eq. (42) is also modified, obtaining
@@
Although not shown here, the equation for the first order time derivative of $h_{q}$ is also modified. In the modified wave equations we can see terms that are directly proportional to first order time derivatives. These are typically associated to dissipative processes (although this need not always be the case). In this case, the dissipative constant is controlled both by $V_{+}=\zeta+4\eta/3+\rho(\tau_{\mathrm{e}}-\tau_{\mathrm{p}})$, as well as the shear viscosity directly, $\eta$. The transport coefficient $\tau_{\mathcal{Q}}$ does not appear altogether. The bulk viscosity only appears through $V_{+}$, and the remaining transport coefficients only appear through $\tau_{-}$. This suggests the presence of significant correlations between the transport coefficients, when looking at the perturbative relaxation of viscous NSs.
@@ 6.3 Viscous Regime
The remaining equations, on the other hand, are very lengthy and unilluminating.We provide the expressions for the remaining equations, as well as those for the odd parity sector, via a Wolfram language package Redondo-Yuste (2024). We provide these equations in a way that is ready for a numerical implementation, either in the time or in the frequency domain. The only things left to be specified are the dependence of the transport coefficients on the fundamental properties of the star, e.g., $\eta=\eta(\mathrm{e})$, and the equation of state, $\mathrm{p}=\mathrm{p}(\mathrm{e})$.
@@ 7 Conclusions
We have employed the formalism of Gerlach and Sengupta Gerlach and Sengupta (1979a, b, 1980), as applied by Gundlach and Martín-García Martin-Garcia and Gundlach (1999); Gundlach and Martin-Garcia (2000); Martin-Garcia and Gundlach (2001) to reduce the perturbed Einstein equations to a set of scalar equations, in terms of gauge invariant variables. We have focused our attention on radiative multipoles, i.e., multipoles with angular index $l\geq 2$, leaving the study of the lower multipoles (including radial perturbations) to future work.
@@
In the polar sector, the situation is more complicated. We have first checked that we recover the perfect fluid results in the appropriate limit. Then, we show how these are modified in the ICA, i.e., freezing the fluid degrees of freedom. In this case, there are two wave-like equations for gravitational wave modes, as well as an elliptic equation, which acts as a constraint. Remarkably, the equations are modified through the inclusion of dissipative terms, i.e., terms proportional to first order time derivatives, proportionally only to two particular combinations of the transport coefficients. This shows that although the dimensionality of the parameter space of transport coefficient is large, most of the dynamics is perhaps captured by some simple combinations of transport coefficients.
Finally, we provide a Wolfram language package Redondo-Yuste (2024) and scripts with the perturbative equations including all dissipative effects, also in the polar sector. We studied their causal structure. We show the existence of a single propagating degree of freedom, associated to even parity GW modes, as well as three fluid modes. One of these fluid modes has the same propagation speed as the novel fluid mode in the axial sector, which we dub a shear mode. This mode does not have a perfect fluid counterpart. We argue that the last pair of modes is composed by the usual fluid mode in the even parity sector (associated to density fluctuations), whose propagation speed is modified in the presence of viscosity, and another mode which is purely viscous. Remarkably, we are able to recover the causality constraints of the BDNK theory from the analysis of the principal part of these equations, therefore checking the validity of our results.
We did not study the effect of non radiating multipoles ($l=0,1$), as well as the junction conditions that match these perturbations to the perturbations in the exterior of the star (described by the Regge–Wheeler and Zerilli equations, respectively). Moreover, we simplified the problem by assuming that the thermal conductivity was negligible. These are all directions in which we plan to extend our work in the future, which also include an extension to capture the effect of magnetic fields Armas and Camilloni (2022). More interestingly, the equations provided here allow us to improve our understanding of the interaction between GWs and matter. In a companion work Boyanov et al. (2024) we have shown how this occurs for axial perturbations. Studying scattering properties, as well as the spectrum of compact stars under even parity perturbations, using the equations provided in this work, is a natural continuation of this.
@@ Acknowledgements
I am indebted to the developers and maintainers of xAct, which made possible most of this work Mart´ın-Garc´ıa (2008).
@@
I acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184.
@@ Appendix A Gauge Invariance of the perturbations
where we remind the reader that $D_{a}$ is the covariant derivative on the 22–sphere, $\epsilon_{a}^{b}$ is the totally antisymmetric tensor in the two dimensional Lorentzian manifold, and $\mathcal{Y}^{lm}$ are the spherical harmonics, defined as the eigenfunctions of the equation
@@
An arbitrary vector field can also be decomposed in terms of a $2$ dimensional vector $\xi_{A}$, and two scalars $\xi,\chi$, as
@@ Appendix C Coefficients of the odd parity equations
$\mathsf{a}=$ $-\frac{e^{\Phi/2}}{r}\Biggl{(}\frac{\partial\log\eta}{\partial r_{\star}}+\frac{1}{2}\frac{d\Phi}{dr_{\star}}\Biggr{)}\,,\hskip 25.0pt\,$, (C.1)
@@
$\hskip 85.35826pt+re^{(\Phi+\Lambda)/2}\Bigl{(}\frac{d\Lambda}{dr_{\star}}+3\frac{d\Phi}{dr_{\star}}\Bigr{)}\Biggr{]}\,$,
@@
$\mathsf{c}_{2}=$ $e^{-\Lambda/2}\Biggl{[}\tau+4\pi r\Biggl{(}re^{\Lambda}\frac{d\Phi}{dr_{\star}}-6e^{(\Phi+\Lambda)/2}\Biggr{)}\frac{\eta}{\mathcal{K}}-8\pi r^{2}e^{\Lambda}\frac{\eta}{\mathcal{K}}\frac{\partial\log\eta}{\partial r_{\star}}\Biggr{]}\,$,
@@
$\mathcal{K}=\frac{d\Lambda}{dr_{\star}}+\frac{d\Phi}{dr_{\star}}=8\pi re^{(\Phi+\Lambda)/2}(\mathrm{e}+\mathrm{p})\,.$ (C.2)
@@
$\frac{\partial\beta}{\partial t}=-e^{-\Phi/2}\Biggl{(}r\frac{\partial\psi}{\partial r_{\star}}+e^{(\Phi-\Lambda)/2}\psi\Biggr{)}\,$, (C.3)